A machine produces parts with lengths that are normally distributed with σ = 0.6. A sample of 19 parts has a mean length of 75.28.

(a) Give a point estimate for μ. (Give your answer correct to two decimal places.)

(b) Find the 98% confidence maximum error of estimate for μ. (Give your answer correct to three decimal places.)

(c) Find the 98% confidence interval for μ. (Give your answer correct to three decimal places.)

Lower Limit Upper Limit You may need to use the appropriate table in Appendix B to answer this question.
Answer #1

Given,

sigma= 0.6

sample mean = ar{x} = 75.28

Sample size = n = 19

a )

Sample mean ( ar{x} ) is point estimate of population mean mu

point estimate for μ = ar{x} = 75.28

b)

Maximum error of estimate for μ is ,

ME = z_{alpha /2}*rac{sigma }{sqrt{n}}

Where, z_{alpha /2} is critical value at given confidence level

Confidence level = 98% = 0.98

Significance level = alpha = 1 - 0.98 = 0.02 , alpha/2 = 0.01

z_{alpha /2} = 0.01 = 2.3264                    { Using Excel, = NORMSINV( 0.01 ) = 2.3264 }

So, 98% confidence maximum error of estimate for μ is,

0.6 ME = 2.3264 *

ME = 0.320228 approx0.320

c)

98% confidence interval for μ is,

Lower Limit = ar{x} - ME = 75.28 - 0.320 = 74.960

Upper Limit = ar{x} + ME = 75.28 + 0.320 = 75.600